You have a child who will start college in 10 years, and you plan to set aside $1,000 a year for her college education during that period. You estimate that you will earn an annual interest rate of 7% on your investment. What amount can you expect to have available for your child when they start college?
Answer #1

[(1+r) -1 FV of Annuity = P P= Periodic Payment r=rate per period n = number of periods

FV = 1000 * (1+0.07) 10 -1 0.07

FV = 1000 * 13.81645

FV = $13,816.45